### v1

# Deep-Relative-Trust-Based Diffusion for Decentralized Deep Learning

Abstract
Decentralized learning strategies allow a collection of agents to learn efficiently from local data sets without the need for central aggregation or orchestration. Current decentralized learning paradigms typically rely on an averaging mechanism to encourage agreement in the parameter space. We argue that in the context of deep neural networks, which are often over-parameterized, encouraging consensus of the neural network outputs, as opposed to their parameters can be more appropriate. This motivates the development of a new decentralized learning algorithm, termed DRT diffusion, based on deep relative trust (DRT), a recently introduced similarity measure for neural networks. We provide convergence analysis for the proposed strategy, and numerically establish its benefit to generalization, especially with sparse topologies, in an image classification task.
Index Terms:
Decentralized learning, distributed learning, deep neural networks, deep relative trust.
I Introduction
We consider a collection of $K$ agents, each equipped with a neural network of identical architecture consisting of $L$ layers. We denote the parameters of the $p$-th layer of agent $k$ by $w_{k}^{(p)}$, and collect for brevity $w_{k}\triangleq\mathrm{col}\{w_{k}^{(p)}\}_{p=1}^{P}$. We denote the output of the neural network with parameterization $w_{k}$ to the input $x$ by $f(x;w_{k})$. To train the parameters of the network to fit a random pair of inputs $\bm{x}$ and outputs $\bm{y}$, we define a loss function $Q(w_{k};\bm{x},\bm{y})$, giving rise to the risk:
$J_{k}(w_{k})=\mathds{E}Q(w_{k};\bm{x},\bm{y})$ (1)
Most algorithms for decentralized learning pursue solutions to the consensus optimization problem [1, 2, 3, 4]:
$\min_{w}J(w)\quad\text{where }J(w)\triangleq\frac{1}{K}\sum_{k=1}^{K}J_{k}(w)$ (2)
For the purposes of motivating this work, it will be sufficient to review the diffusion algorithm [2]:
$\bm{\psi}_{k,i}$ $=\bm{w}_{k,i-1}-\mu\widehat{\nabla J}_{k}\left(\bm{w}_{k,i-1}\right)$ (3a)
$\bm{w}_{k,i}$ $=\sum_{\ell=1}^{K}a_{\ell k}\bm{\psi}_{\ell,i}$ (3b)
Here, $\widehat{\nabla J}_{k}(\cdot)$ denotes a stochastic approximation of the true local gradient $\nabla J_{k}(\cdot)$ based on data available at time $i$. A common choice is $\widehat{\nabla J}_{k}(\cdot)\triangleq\nabla Q_{k}(\cdot;\bm{x}_{i},\bm{y}_{i})$, although many variants such as mini-batch approximations are possible. The scalars $a_{lk}$ are combination weights satisfying:
$a_{\ell k}\geq 0,\quad\sum_{\ell\in\mathcal{N}_{k}}a_{\ell k}=1,\quad a_{\ell k}=0\text{ if }\ell\notin\mathcal{N}_{k}$ (4)
where $\mathcal{N}_{k}$ denotes the neighbourhood of agent $k$. Step (3a) corresponds to a local gradient descent step, encouraging (local) optimality of the intermediate weight vectors $\bm{\psi}_{k,i}$, while (3b) corresponds to a consensus averaging step, driving local weight vectors towards a common solution. Convergence of (3a)–(3b) to a minimizer, first- or second order-stationary point of [2] has been established under various conditions on the local objectives $J_{k}(\cdot)$ [1, 5, 3, 6]. In all of these cases, an important quantity in the rate of convergence is the mixing rate of $A=[a_{\ell k}]$, given by its second largest eigenvalue $\lambda_{2}$. As discussed in [7], an effective albeit suboptimal method for achieving fast convergence is the Metropolis rule:
$a_{\ell k}=\begin{cases}1/\max\left\{n_{k},n_{\ell}\right\},&\text{ if }k\neq\ell\text{ and }\ell\in\mathcal{N}_{k}\\ 1-\sum_{\ell\in\mathcal{N}_{k}\backslash\{k\}}a_{\ell k},&\text{ when }k=\ell\\ 0,&\text{ otherwise }\end{cases}$ (5)
where $n_{k}\triangleq|\mathcal{N}_{k}|$ is the degree of node $k$.
Observe that the combination step (3b) drives the parameters $\bm{w}_{k,i}$ towards consensus. Other strategies for decentralized optimization and learning, such as EXTRA [8] or Exact diffusion [9] rely on primal-dual constructions, while gradient-tracking based algorithms such as NEXT [10] or DIGing [11] rely on the dynamic consensus mechanism. All of them, however, in a manner analogous to (3b) encourage equality in the parameter space.
In this work, we deviate from this paradigm and argue that when training neural networks, it may be more appropriate to encourage consensus in the function space. This is because modern neural networks are highly over-parameterized, and as a result the same function may be obtained with highly variable parameterizations. Indeed, some evidence exists in the literature that allowing for some variability in local parameterizations during training can aid generalization performance of deep neural networks [12]. Motivated by these observations, we formulate a penalty-based variant of (2), where agents are allowed to maintain distinct local parameterizations, provided that the resulting function is approximately the same. We provide analytical convergence guarantees for the resulting algorithm, and demonstrate numerically its benefit in deep learning.
II DRT Diffusion
It can be verified that the diffusion algorithm (3a)–(3b) is equivalent to incremental stochastic gradient descent applied to the following penalized variant of (2[9]:
$\min_{w_{k}}J_{k}(w_{k})+\frac{\eta}{2}\sum_{\ell\in\mathcal{N}_{k}}c_{\ell k}{\lVert w_{k}-w_{\ell}\lVert^{2}}$ (6)
where $c_{\ell k}>0$ if and only if $\ell\in\mathcal{N}_{k}$. We can observe that under this relaxation, individual agents are allowed to maintain local models $w_{k}$, but are nevertheless encouraged to ensure $w_{k}\approx w_{\ell}$ through the penalty term $\frac{\eta}{2}\sum_{\ell\in\mathcal{N}_{k}}c_{\ell k}{\lVert w_{k}-w_{\ell}\lVert^{2}}$. In this work, instead of penalizing difference between parameters, we propose to penalize the difference between neural network outputs:
$\min_{w_{k}}J_{k}(w_{k})+\frac{\eta}{2}\sum_{\ell\in\mathcal{N}_{k}}c_{\ell k}\frac{\lVert f(x;w_{k})-f(x;w_{\ell})\lVert^{2}}{\lVert f(x;w_{\ell})\lVert^{2}}$ (7)
One challenge of this formulation is that differentiating the penalty term $\frac{\eta}{2}\sum_{\ell\in\mathcal{N}_{k}}c_{\ell k}\frac{\lVert f(x;w_{k})-f(x;w_{\ell})\lVert^{2}}{\lVert f(x;w_{\ell})\lVert^{2}}$ is computationally prohibitive. Instead, we will replace this penalty by an upper bound, motivated by the Deep Relative Trust (DRT) distance measure [13]:
$\frac{\lVert f(x;w_{\ell})-f(x;w_{k})\lVert}{\lVert f(x;w_{k})\lVert}\leq\prod_{p=1}^{L}\left(1+\frac{\lVert w_{k}^{(p)}-w_{\ell}^{(p)}\lVert}{\lVert w_{k}^{(p)}\lVert}\right)-1$ (8)
Using arguments analogous to those in [13], we can establish a quadratic variant:
$\frac{\lVert f(x;w_{k})-f(x;w_{\ell})\lVert^{2}}{\lVert f(x;w_{\ell})\lVert^{2}}$ (9)
$\leq$ $2^{L+1}\prod_{p=1}^{L}\left(1+\frac{\lVert w_{k}^{(p)}-w_{\ell}^{(p)}\lVert^{2}}{\lVert w_{\ell}^{(p)}\lVert^{2}}\right)+2$
which will be more conveniently differentiated than (8). This gives rise to the penalized optimization problem:
$\min_{w_{k}}J_{k}(w_{k})$ (10)
$+$ $\frac{\eta}{2}\sum_{\ell\in\mathcal{N}_{k}}c_{\ell k}\left[2^{L+1}\prod_{p=1}^{L}\left(1+\frac{\lVert w_{k}^{(p)}-w_{\ell}^{(p)}\lVert^{2}}{\lVert w_{\ell}^{(p)}\lVert^{2}+\kappa}\right)+2\right]$
where we added $\kappa\geq 0$ for numerical stability. The parameter $\eta$ controls the penalty associated with differences in neural network outputs, and is chosen as $\eta=\frac{1}{\mu}$, where $\mu$ represents the learning rate for neural network training.
After applying an incremental stochastic gradient construction to (10) and subsequently normalizing the combination weights, we obtain the DRT Diffusion algorithm:
$\bm{\psi}_{k,i}=\bm{w}_{k,i-1}-\mu\widehat{\nabla J}_{k}\left(\bm{w}_{k,i-1}\right)$ (11)
$\bm{w}_{k,i}^{(p)}=\sum_{\ell\in\mathcal{N}_{k}}\bm{a}_{\ell k,i}^{(p)}\bm{\psi}_{\ell,i}^{(p)}$
Observe that the resulting algorithm resembles the classical diffusion strategy (3a)–(3b), except that the combination weights as a result of the penalty in (10) now also depend on the layer index $p$ and time $i$. More specifically, the combination weights $\bm{a}_{\ell k,i}^{(p)}$ are given by:
$\bm{a}_{\ell k,i}^{(p)}=\frac{\tilde{\bm{a}}_{\ell k,i}^{(p)}}{\sum_{\ell\in\mathcal{N}_{k}}\tilde{\bm{a}}_{\ell k,i}^{(p)}}$ (12)
where
$\tilde{\bm{a}}_{\ell k,i}^{(p^{*})}=\begin{cases}\min\left(\accentset{\approx}{\bm{a}}_{\ell k,i}^{(p^{*})},N\min_{\ell}\left(\accentset{\approx}{\bm{a}}_{\ell k,i}^{(p^{*})+}\right)\right)&\ell\neq k\\ \frac{c_{kk}}{n_{k}-1}\sum_{\ell\in\mathcal{N}_{k}\backslash\{k\}}\tilde{\bm{a}}_{\ell k,i}^{(p^{*})}&\ell=k\end{cases}$ (13)
$\accentset{\approx}{\bm{a}}_{\ell k,i}^{(p^{*})}=\begin{cases}c_{\ell k}\frac{2^{L+1}\prod_{p=1}^{P}(1+\frac{\lVert\bm{w}_{k,i}^{(p)}-\bm{w}_{\ell,i}^{(p)}\lVert^{2}}{\lVert\bm{w}_{\ell,i}^{(p)}\lVert^{2}+\kappa})}{\lVert\bm{w}_{\ell,i}^{(p^{*})}\lVert^{2}+\lVert\bm{w}_{k,i}^{(p^{*})}-\bm{w}_{\ell,i}^{(p^{*})}\lVert^{2}+\kappa}&\ell\neq k\\ 0&\ell=k\end{cases}$ (14)
Here, $n_{k}$ represents the number of neighbors associated with agent $k$. The parameter $N$, which satisfies $N\geq 1$, determines the smallest possible positive entry in the mixing matrices. This minimum entry is given by $\frac{1}{(K-1)N+1}$, where $K$ denotes the total number of agents. Furthermore, $\min_{\ell}\left(\accentset{\approx}{\bm{a}}_{\ell k,i}^{(p^{*})+}\right)$ refers to the smallest positive value of $\accentset{\approx}{\bm{a}}_{\ell k,i}^{(p^{*})}$ among all indices $\ell$. It can be verified that under this construction:
$\bm{a}_{\ell k,i}^{(p)}\geq 0,\quad\sum_{\ell\in\mathcal{N}_{k}}\bm{a}_{\ell k,i}^{(p)}=1,\quad\bm{a}_{\ell k,i}^{(p)}=0\text{ if }\ell\notin\mathcal{N}_{k}$ (15)
The effect of these combination weights is that parameter deviations are penalized in relation to their effect on the output of the neural network.
III Convergence Analysis
We first need to relate the connectivity of the underlying graph topology captured in $C=[c_{\ell k}]$ to the connectivity of the induced combination matrices $\bm{A}_{i}^{(p)}\triangleq\left[\bm{a}_{\ell k}\right]_{i}^{(p)}$.
Assumption 1 (Strongly-connected graph).
The graph described by the weight matrix $C=[c_{\ell k}]$ is strongly-connected, implying that $C$ is a primitive matrix. ∎
Lemma 1 (Primitive Mixing Matrices).
Under Assumption 1 and the construction (14), the graph represented by the weighted combination matrix $\bm{A}_{i}^{(p)}\triangleq\left[\bm{a}_{\ell k}\right]_{i}^{(p)}$ is primitive for all $p$ and all $i$. ∎
Proof.
Assumption 1 implies that the graph represented by $C$ is primitive. Additionally, Equation (14) indicates that for any $\ell$ and $k$, $\bm{a}_{\ell k,i}^{(p)}=0$ if and only if $c_{\ell k}=0$. Consequently, the graph described by $\bm{A}_{i}^{(p)}$ must also be primitive. ∎
Since we need to account for the stochastic gradient dynamics, the following common assumptions regarding the gradients and their approximations are required.
Assumption 2 (Lipschitz gradients).
For each $k$, the gradient $\nabla J_{k}(\cdot)$ is Lipschitz, namely, for any $x,y\in\mathbb{R}^{M}$:
$\left\|\nabla J_{k}(x)-\nabla J_{k}(y)\right\|\leq\delta\|x-y\|$ (16)
Assumption 3 (Bounded gradients).
For each $k$, the gradient $\nabla J_{k}(\cdot)$ is bounded, namely, for any $x\in\mathbb{R}^{M}$ :
$\left\|\nabla J_{k}(x)\right\|\leq G$ (17)
Assumption 4 (Gradient noise process).
For each $k$, the gradient noise process is defined as
$\bm{s}_{k,i}\left(\bm{w}_{k,i-1}\right)=\widehat{\nabla J}_{k}\left(\bm{w}_{k,i-1}\right)-\nabla J_{k}\left(\bm{w}_{k,i-1}\right)$ (18)
and satisfies
$\mathbb{E}\left[\bm{s}_{k,i}\left(\bm{w}_{k,i-1}\right)\mid\bm{w}_{k,i-1}\right]$ $=0$ (19)
$\mathbb{E}\left[\left\|\bm{s}_{k,i}\left(\bm{w}_{k,i-1}\right)\right\|^{2}\mid\bm{w}_{k,i-1}\right]$ $\leq\sigma_{k}^{2}$
for some non-negative constants $\sigma_{k}^{2}$. ∎
III-A Network Disagreement and Descent
Lemma 2 (Network Disagreement).
Under Assumptions 1-5 and for sufficiently small step-sizes $\mu$, the network disagreement is bounded as:
$\mathbb{E}\sum_{k=1}^{K}[\bm{\phi}_{i}]_{k}\left\|\bm{w}_{k,i}-\bm{w}_{c,i}\right\|^{2}\leq\frac{16\mu^{2}\left(1+K\right)}{\left(1-\bar{\xi}\right)^{2}}\left(G^{2}+\max_{k}\sigma_{k}^{2}\right)$ (20)
The deviation is measured relative to the weighted centroid, defined as:
$\bm{w}_{c,i}\triangleq\sum_{k=1}^{K}[\bm{\phi}_{i}]_{k}\bm{w}_{k,i}$ (21)
where $[\bm{\phi}_{i}]_{k}$ represents the $k$-th entry of the time-varying weight vector $\bm{\phi}_{i}$.
The mixing rate $\bar{\xi}$ is given by:
$\bar{\xi}=\sqrt{1-\frac{\frac{\left(\frac{1}{(K-1)N+1}\right)^{K+2}}{K}}{\mathrm{D}(\mathbb{C})\mathrm{U}(\mathbb{C})\left[1-\left(K-1\right)\frac{\left(\frac{1}{(K-1)N+1}\right)^{K}}{K}\right]^{2}}}$ (22)
with the upper bound:
$\bar{\xi}\leq\sqrt{1-\frac{1}{\left[\left(K-1\right)N+1\right]^{K+2}K\mathrm{D}(\mathbb{C})\mathrm{U}(\mathbb{C})}}<1$ (23)
Here, $\mathrm{D}(\mathbb{C})$ and $\mathrm{U}(\mathbb{C})$ represent the graph diameter and the maximal edge-utility of the graph $\mathbb{C}$, which is characterized by the weight matrix $C$, as described in [14]. ∎
Proof.
The argument combines Lemma 3.3, 3.4 and 4.2 from [14] to account for the time-varying nature of the combination weights $\bm{a}_{\ell k,i}^{(p)}$ with techniques from [15] to establish clustering. It is omitted due to space limitations. ∎
Having established clustering of the agents around a suitably chosen reference point, we can now establish convergence to a first-order stationarity. For simplicity, for this theorem only, we restrict ourselves to the setting of IID data, meaning that agents observe data that is identically distributed. Formally:
Assumption 5 (Common Objective Functions).
The local data distributions are identical, and hence $J_{k}(w)=J_{\ell}(w)$ for all $k$ and $\ell$.
Theorem 1 (Descent Relation).
Under Assumptions 1-5 and for sufficiently small step-sizes, we have:
$\mathbb{E}\left\{J\left(\bm{w}_{c,i}\right)\mid\bm{w}_{c,i-1}\right\}\leq J\left(\bm{w}_{c,i-1}\right)$ (24)
whenever
$\left\|\nabla J\left(\bm{w}_{c,i-1}\right)\right\|^{2}\geq\mu\frac{c_{2}}{c_{1}}$ (25)
where $c_{1}$ and $c_{2}$ are constants indpeendent of $\mu$. ∎
Proof.
The argument is analogous to [6] and omitted due to space limitations. ∎
This theorem ensures the proposed algorithm is a decent recursion for the network centroid in the IID scenario whenever the network centroid $\bm{w}_{c,i}$ is not $O(\mu)$-stationary. Then from Crollollary 1 in [16], we can conclude that at some time $i^{\star}\leq O\left(1/\mu^{2}\right)$, the network centroid will reach an $O(\mu)$-mean-square-stationary point.
IV Simulation Results
IV-A Simulation Setup
Since the convergence guarantees are proven in the IID case, we aim to explore whether DRT diffusion performs well in the non-IID case. For this purpose, we utilise the CIFAR-10 dataset for an image classification task. We deploy 16 agents in a decentralized system, each equipped with a non-IID local dataset sampled from CIFAR-10 without replacement. To ensure the local datasets are non-IID, we first randomly select the number of image classes present at each agent, ranging between 5 and 8. Then, the total number of samples for each agent is randomly chosen to be between 1,500 and 2,000.
To explore the performance of the DRT diffusion algorithm across different network topologies, we experiment with a ring, Erdős–Rényi model with the edge probability of 0.1, and Hypercube.
We compare the performance of the DRT diffusion algorithm with the classical diffusion algorithm [2] using the optimal mixing matrix as described in [7] to ensure rapid convergence. Both algorithms are tested by locally training ResNet-20s for one epoch with a batch size of 128. After local training, the algorithms perform 3 consensus steps. The primary purpose of implementing consecutive consensus steps is to reduce the number of epochs needed to reach a steady state [12]. Without this approach, the simulation would become too computationally expensive. For the mixing matrix construction, $N$ is chosen to be $2K$.
IV-B Results and Discussion
We compare performance based on steady-state performance, convergence rate, and the generalisation gap, which measures the difference between training and test accuracy across all experiments. In particular, we examine the feature of DRT diffusion that facilitates distinct local parameterisations.
TABLE I: Steady-state test accuracy with different topologies
Topology $\lambda_{2}$ Classical diffusion DRT diffusion
Ring 0.949 66.21% 69.60%
Erdős–Rényi 0.905 65.55% 69.22%
Hypercube 0.600 71.05% 72.00%
(a) Ring $\lambda_{2}=0.949$
(b) Erdős–Rényi $\lambda_{2}=0.905$
(c) Hypercube $\lambda_{2}=0.600$
Figure 1: Learning curves for a decentralized network with 16 agents, employing ResNet-20 on CIFAR-10 with non-IID data at each agent
IV-B1 Steady-state Performance
The steady-state test accuracies are displayed in Table I, and the learning curves along with generalisation gaps are depicted in Figures 1 and 2. The results in Table I show that DRT diffusion surpasses classical diffusion in steady-state performance, especially in sparse topologies. However, for the well-connected Hypercube topology, the performance difference is minimal. Interestingly, DRT diffusion exhibits greater fluctuations than classical diffusion in the Hypercube, suggesting it may not be the optimal choice for well-connected topologies. These findings indicate that DRT diffusion, by promoting consensus in the function space, facilitates efficient information exchange among agents.
IV-B2 Convergence Rate
Figure 1 shows that DRT diffusion achieves the same convergence rate as classical diffusion with the fast-mixing Metropolis rule. This is notable since the Metropolis rule is constructed to result in fast information diffusion, while the proposed DRT-based scheme is only designed to ensure equality in neural network outputs.
IV-B3 Generalisation Gap
The generalisation gap indicates how well the model performs on unseen data. Figure 2 demonstrates that DRT diffusion has a smaller generalisation gap compared to classical diffusion with sparse topologies. This suggests that the proposed algorithms has an implicit bias towards minimizers that generalize better, warranting further investingation along the lines of [17, 18].
(a) Ring $\lambda_{2}=0.949$
(b) Erdős–Rényi $\lambda_{2}=0.905$
(c) Hypercube $\lambda_{2}=0.600$
Figure 2: Generalisation gap for a decentralized network with 16 agents, employing ResNet-20 on CIFAR-10 with non-IID data at each agent
V Conclusion
This work introduces the DRT diffusion method, which promotes consensus in the function space. Analytical convergence guarantees are established in the IID scenario, while simulations on CIFAR-10 with locally non-IID data distribution empirically establish improved steady-state performance over classical diffusion with sparse topologies, particularly in reducing the generalization gap.